Graph this equation by plotting three points: y = 3x - 2.


What are your points?

Where does your line cross the x-axis(write as an ordered pair, or if it is not a "nice" point, provide the two numbers it lies between)?

Where does your line cross the y-axis(write as an ordered pair)?

Is the origin on the right or left side of the line?
1
Expert's answer
2020-02-27T11:28:05-0500

Solution:

y = 3x - 2 --------- (1)


(1)

Put x= 0 we get y = "-" 2 , so the first point is (0,"-" 2).


Put y= 0 we get x= "\frac{2}{3}" , so the second point is ( "\frac{2}{3}" , 0).


Put x=1 we get y = 1 , so the third point is (1, 1).


Graph is as follows:






(2)

Points are (0,-2) , ("\frac{2}{3}" ,0) ,(1,1)


(3)

Line crosses the x-axis at the point ("\frac{2}{3}" , 0) // the y-coordinate (ordinate) of the x-axis is always zero.


(4)

Line crosses the y-axis at the point (0,-2) // the x-coordinate (abscissa) of the y-axis is always zero.


(5)

Origin(0,0) lies on the left hand side of the line.